# Approximate Degrees of Multisymmetric Properties with Application to Quantum Claw Detection

### v1

@@ 1.3.2 Application to Claw Detection
$F\leq G\leq F^{2}$ $O\left(\min\left\{(FG)^{1/3},G^{1/2+\epsilon}M^{1/6}\right\}\right)$ [30, 7] $O((FG)^{1/3})$ [30]
@@
We use Theorem 1 to derive a lower bound for the claw detection with a small range. For this, we need a lower bound on the minimum degree of a polynomial that approximates the claw property (for input functions with a possibly large common range). It turns out that for any constant $20<\epsilon<1/$ 2, the minimum degree of a polynomial that $\epsilon$-approximates the claw property for function pairs $f\colon[F]\to[FG]$ and $g\colon[G]\to[FG]$ is $\Omega\left((FG)^{1/3}\right)$. By combining this with Theorem 1, we can obtain a lower bound on the minimum approximate degree of the claw property for the range size at least $F+G$ of input functions.
@@
Let $F,G\in\mathbb{N}$ be such that $F\leq G$. For any constant $20<\epsilon<1/$ 2 and any $M\geq F+G$, the minimum degree of a polynomial that $\epsilon$-approximates the claw property for a function pair in $[M]^{[F]}\times[M]^{[G]}$ is $\Theta(\sqrt{G}+(FG)^{1/3})$, where the lower bound $\Omega(\sqrt{G})$ holds for all $M\geq 2$.
By the polynomial method, Theorem 3 implies the optimal lower bound on the bounded-error quantum query complexity of both the claw detection and the claw finding, where the optimality follows from the quantum algorithm [30] that finds a claw.
@@
Let $F,G\in\mathbb{N}$ be such that $F\leq G$. For every $M\geq F+G$, the (query-)optimal quantum algorithm that computes a claw (or detects the existence of a claw) with a constant error probability requires $\Theta(\sqrt{G}+(FG)^{1/3})$ queries for a given pair of functions in $[M]^{[F]}\times[M]^{[G]}$. More precisely, the optimal quantum query complexity is $\Theta((FG)^{1/3})$ for every $M\geq F+G$ if $F\leq G\leq F^{2}$ and $\Theta(\sqrt{G})$ for every $M\geq 2$ if $G>F^{2}$.
Ambainis, Balodis and Iraids [7] investigated upper and lower bounds for the claw problem for input function pairs with much smaller ranges. They proved the lower bound $\Omega\left(N^{1/2}M^{1/6}\right)$ when the two input functions has the same domain size $N$ and the range size $M$ with $2\leq M<N$. Their proof is based on the lower bound $\Omega(N^{2/3})$ on the number of quantum queries required to detect a claw for a given pair of functions in $[N]^{[N]}\times[N]^{[N]}$, which follows by reducing from the element distinctness. However, if the domains of two input functions are different in size, their proof breaks down. By generalizing their proof using Theorem 4, we obtain a lower bound for functions with different-size domains and very small common ranges.
@@
Let $F,G,M\in\mathbb{N}$ be such that $F\leq G\leq F^{2}$ and $M<F+G$. Then, for every $M\in[2,F+G-1]$, the quantum query complexity of detecting the existence of a claw for a given function pair $(f,g)$ in $[M]^{[F]}\times[M]^{[G]}$ is lower-bounded by
@@
Note that the condition of $F\leq G\leq F^{2}$ and $M<F+G$ imposed in Theorem 5 means the parameter range that Theorem 4 does not cover. Setting $F=G=N$ in the theorem recovers the lower bound $\Omega\left(N^{1/2}M^{1/6}\right)$ [7].
To evaluate our bound in Theorem 5, we will compare it with the lower bound obtained by straightforwardly reducing the claw detection for two functions with domains of the same size and applying the lower bound given in  [7]. Let us write $\mathsf{claw}_{(a,b)\to c}$ to mean the decision problem of detecting the existence of a claw for given two functions in $[c]^{[a]}\times[c]^{[b]}$. The reduction from $\mathsf{claw}_{(G,G)\to M}$ to $\mathsf{claw}_{(F,G)\to M}$ is as follows, where we assume for simplicity that $G$ is divisible by $F$, i.e., $F|G$ (it is easy to extend the reduction to the general case by padding dummy elements to one of the domains). Let $(f,g)$ be a pair of input functions to $\mathsf{claw}_{(G,G)\to M}$. We first partition the domain $[G]$ of $f$ into $G/F$ blocks of size $F$: $B_{1},\dots,B_{G/F}$. For randomly chosen $i\in[G/F]$, we run a $Q$-query quantum algorithm for $\mathsf{claw}_{(F,G)\to M}$ on $f$ restricted to the domain $B_{i}$ and $g$, and apply the quantum amplitude amplification [12, 19] to the whole algorithm. The resulting quantum algorithm makes $O\left(\sqrt{G/F}\cdot Q\right)$ quantum queries. Since the quantum query complexity of $\mathsf{claw}_{(G,G)\to M}$ for $M<G$ is $\Omega(G^{1/2}M^{1/6})$ [7], we have $Q=\Omega(F^{1/2}M^{1/6})$. Together with trivial lower bound $\Omega(\sqrt{G})$, we obtain the lower bound $\Omega(\sqrt{G}+F^{1/2}M^{1/6})$. This bound is weaker than our $\Omega(\sqrt{G}+F^{1/3}G^{1/6}M^{1/6})$. To see this, assume $G=F^{1+\epsilon}$ for $0<\epsilon\leq 1$, satisfying $F<G\leq F^{2}$. In this case, we have
@@
Tab. missing 1 summarizes the quantum query complexity for claw detection for functions $f\colon[F]\to[M]$ and $g\colon[G]\to[M]$, assuming $F\leq G$ without loss of generality. The upper bound $O(G^{1/2+\epsilon}M^{1/4})$ in the case of $M<F+G$ for $F\leq G\leq F^{2}$ follows from the upper bound [7] for $\mathsf{claw}_{(G,G)\to M}$ (i.e., the claw detection for the same domain sizes) via a simple embedding the domain $[F]$ into $[G]$.
The most interesting open problem is obviously to find the tight bound in the case of $M\leq F+G$ for $F\leq G\leq F^{2}$. More precisely, it follows from Tab. missing 1 that the tight bound $\Theta((FG)^{1/3})$ for $M\geq F+G$ also holds for all $M=\Omega(G)$. On the other hand, if $M=\Theta(1)$, then $\Omega(\sqrt{G})$ is tight, since Grover’s search algorithm achieves $O(\sqrt{G})$ queries. Thus, the tight bound is open when $M$ is asymptotically smaller than $G$ and larger than $2$.
@@ 1.4 Technical Outline
The starting point of the proofs of Theorems 1 and 2 is the proof idea for the single-symmetric properties introduced by Ambainis [4] (we hereafter refer to "single-symmetric" just as "symmetric"). Let $\Phi$ and $\Phi^{\prime}$ be the symmetric Boolean functions defined in Sec. missing  1.2. Recall that $\Phi$ and $\Phi^{\prime}$ are functions in $\{x_{ij}\colon(i,j)\in[n]\times[m]\}$ and $\{x_{ij}\colon(i,j)\in[n]\times[n]\}$, respectively. In addition, define $z_{j}$ as the number of preimages of $j$ via $f$, i.e., $z_{j}=\left|f^{-1}\left(j\right)\right|$. Then, Ambainis’s proof consists of two claims. The first claim is that the following are equivalent:
@@
These claims together imply that the minimum degree of a polynomial that approximates $\Phi$ is at most $d$, if and only if the minimum degree of a polynomial that approximates $\Phi^{\prime}$ is at most $d$.
We want to generalize these two claims to the bisymmetric case (the $k$-symmetric case is a straightforward extension of the bisymmetric case). Recall in the bisymmetric case that $f\colon[F]\to[M]$ is identified with $FM$ Boolean variables $\{x_{ij}:i\in\left[F\right],j\in\left[M\right]\}$, and $g\colon[G]\to[M]$ is identified with $GM$ Boolean variables $\{y_{kj}:k\in\left[F\right],j\in\left[M\right]\}$. In adition, we define $z_{j}$ as the number of preimages of $j$ via $f$, i.e., $z_{j}=\left|f^{-1}\left(j\right)\right|$, and define $w_{j}$ as the number of preimages of $j$ via $g$, i.e., $w_{j}=\left|g^{-1}\left(j\right)\right|$. Let $\Phi$ and $\Phi^{\prime}$ be the bisymmetric Boolean function in $\{x_{ij}\}$ and $\{y_{kj}\}$ as defined in Sec. missing  1.3. The first claim in the bisymmetric case is that the following are equivalent:
@@
For the first claim in the symmetric case, it is easy to show that item $1$ implies item $2$ by using the trivial relation $z_{j}=\sum_{i}x_{ij}$. This works also in the bisymmetric case. To show the converse in the symmetric case, we consider the average $\mathsf{E}_{\pi}[P]$ of $P$ over all permutations $\pi$ over the domains of $f$, since $\mathsf{E}_{\pi}[P]$ also approximates $\Phi$ due to the symmetry of $\Phi$ over the domains. Since we are interested in the degree of the resulting polynomial, it suffices to examine the average of each product term $x_{i_{1}j_{1}}\cdots x_{i_{k}j_{k}}$ in $P$ and it turns out that $\mathsf{E}_{\pi}\left[x_{i_{1}j_{1}}\cdots x_{i_{k}j_{k}}\right]$ is a polynomial in $z_{j}$’s of degree at most $d$. In the bisymmetric case, we take average $\mathsf{E}_{\pi_{F}\times\pi_{G}}[P]$ of $P$ over all permutations $\pi_{F}$ over the domains of $f$ and all permutations $\pi_{G}$ over the domains of $g$. Note that each product term in $P$ is of the form of $x_{i_{1}j_{1}}\cdots x_{i_{m}j_{m}}y_{k_{1}j_{1}}\cdots y_{k_{l}j_{l}}$. To obtain the average of each term, we apply permutations $\pi_{F}$ and $\pi_{G}$ independently and thus it holds that
@@
For the second claim, it is more involved to generalize the symmetric case to the bisymmetric case since the permutation over the range is common to the two functions. It is easy to see that item (i) implies (ii) in the symmetric case since we can obtain $Q^{\prime}$ by setting $z_{n+1}=\cdots=z_{m}=0$ in $Q$. This works also in the bisymmetric case. For the converse, let us first review the symmetric case. Since every $z:=(z_{1},\dots,z_{m})$ corresponding to a function $f:\left[n\right]\rightarrow\left[m\right]$ has at most $n$ nonzero coordinates, there exists a permutation $\sigma$ over $[m]$ that moves all non-zero coordinates to the first $n$ coordinates, that is, $\sigma z=\left(z_{\sigma^{-1}\left(1\right)},\ldots,z_{\sigma^{-1}\left(m\right)}\right)$ satisfies that $z_{\sigma^{-1}\left(j\right)}=0$ for every $\sigma^{-1}\left(j\right)\geq n+1$. This action of $\sigma$ does not change the value of $\Phi$: $\Phi\left(f\right)=\Phi\left(\sigma f\right)$. Since $\sigma f$ can be regarded as a function in $\left[n\right]^{\left[n\right]}$, it must hold that $\Phi\left(\sigma f\right)=\Phi^{\prime}\left(\sigma f\right)$. By the assumption, $\Phi^{\prime}\left(\sigma f\right)$ is approximated by $Q^{\prime}\left(\sigma z\mid_{n}\right)$, where $\sigma z\mid_{n}$ is the first $n$ variables of $\sigma z$, i.e., $z_{\sigma^{-1}\left(1\right)},\ldots,z_{\sigma^{-1}\left(n\right)}$. We then define from $Q^{\prime}$ a symmetric polynomial $Q$ in $m$ variables, $z_{1},\ldots,z_{m}$, such that, for $z$ corresponding to an arbitrary $f$,
@@
where $\sigma$ is a particular permutation defined above depending on $f$. This implies that $Q\left(z\right)$ approxiamtes $\Phi\left(f\right)$, since $Q^{\prime}\left(\sigma z\mid_{n}\right)$ approximates $\Phi^{\prime}\left(\sigma f\right)=\Phi\left(\sigma f\right)=\Phi\left(f\right)$. Since we can assume without loss of generality that $Q^{\prime}$ is a symmetric polynomial, the fundamental theorem of symmetric polynomials implies that $Q^{\prime}$ is represented as a polynomial in the elementary polynomials over $z_{1},\dots,z_{n}$. We then replace these elementary polynomials in $Q^{\prime}$ with the corresponding polynomials over $z_{1},\dots,z_{m}$ to obtain $Q$.
In the bisymmetric case, $Q$ is a polynomial in two sets of variables: $z_{1},\dots,z_{M}$ and $w_{1},\dots,w_{M}$; $Q^{\prime}$ is a polynomial in two sets of variables $z_{1},\dots,z_{M^{\prime}}$ and $w_{1},\dots,w_{M^{\prime}}$ such that $F+G\leq M^{\prime}<M$, since there exists a permutation $\sigma$ over $[M]$ that moves all non-zero coordinates in $z_{1},\dots,z_{M}$ and $w_{1},\dots,w_{M}$ to the first $M^{\prime}$ coordinates, respectively. To obtain $Q$ from $Q^{\prime}$, we utilize a certain symmetry of $Q^{\prime}$. However, $Q^{\prime}$ is not invariant under the action of permutations over $[2M^{\prime}]$ to $2M^{\prime}$ variables $z_{1},\dots,z_{M^{\prime}},w_{1},\dots,w_{M^{\prime}}$, but it is invariant under the action of $\tau\times\tau$ for permutation $\tau$ over $[M^{\prime}]$ such that $(z_{1},\dots,z_{M})\mapsto\left(z_{\tau^{-1}\left(1\right)},\ldots,z_{\tau^{-1}\left(M^{\prime}\right)}\right)$ and $(w_{1},\dots,w_{M})\mapsto\left(w_{\tau^{-1}\left(1\right)},\ldots,w_{\tau^{-1}\left(M^{\prime}\right)}\right)$. Here, the notion of multisymmetric polynomials comes in. Roughly speaking, multisymmetric polynomials are those invariant under the action of permutations to the set of vectors of variables, instead of variables. In our case, let us consider the set of vectors $(z_{1},w_{1}),\dots,(z_{M^{\prime}},w_{M^{\prime}})$, each of which consists of two variables. Then, $Q^{\prime}$ is invariant under the action of permutation $\tau$ to this set of vectors. Now that we enter the world of multisymmetric polynomials, we can use the fundamental theorem of those polynomials. More concretely, $Q^{\prime}$ is represented as a polynomial in power-sum polynomials (defined in Sec. missing  2.4) in $(z_{1},w_{1}),\dots,(z_{M^{\prime}},w_{M^{\prime}})$. Then, we can obtain $Q$ by replacing each occurrence of power-sum polynomials in $Q^{\prime}$ to the corresponding polynomials in $(z_{1},w_{1}),\dots,(z_{M},w_{M})$. This approach is easily generalized to the $k$-symmetric case. We just consider the symmetry under the action of permutations to the set of vectors $\left(z_{1}^{(1)},\dots,z_{1}^{(k)}\right),\dots,\left(z_{M^{\prime}}^{(1)},\dots,z_{M^{\prime}}^{(k)}\right)$, where $z_{i}^{(\ell)}=\left\lvert f_{\ell}(j)\right\rvert$.
@@
To provide a lower bound for the claw problem with a small range using Theorem 1, we need a lower bound on the minimum degree of a polynomial that approximates the claw property for some common range size of input functions. For this, we use the Zhang’s reduction [32] from the collision problem to the claw problem. Suppose that there exists a polynomial $P^{\prime}$ of degree $d$ that 0.1-approximates the claw property for functions $f\in[FG]^{[F]}$ and $g\in[FG]^{[G]}$ such that $F\leq G$. Based on $P^{\prime}$, we will then construct a polynomial $P$ of degree $d$ that (41/86)-approximates the collision property for functions $h\in[M]^{[M]}$ for $M=FG$ with the promise that $h$ is either one-to-one or two-to-one. Here we mean by the collision property the Boolean function $\Phi$ which is $\mathsf{true}$ if and only if the input function $h$ is two-to-one. Since the minimum degree of the latter polynomial is $\Omega(M^{1/3})$ [2, 4], we obtain $d=\Omega((FG)^{1/3})$. Theorem 1 with this bound implies Theorem 3, from which we obtain Theorem 4 with the polynomial method.
To obtain polynomial $P$ from $P^{\prime}$, let $P^{\prime}_{ST}$ for disjoint subsets $S,T$ of domain $[M]$ of $h$ be a polynomial of degree $d$ (in variables $x_{ij}$ for $i\in S\sqcup T$ and $j\in[M]$) that $\epsilon$-approximates the claw property on function pairs $[M]^{S}\times[M]^{T}$ to $\{0,1\}$. Then, define a polynomial $P$ as the average of $P^{\prime}_{ST}$ over uniformly random disjoint subsets $S,T$ of size $F$ and $G$, respectively:

### v2

@@ 1.3.2 Application to Claw Detection
$F\leq G\leq F^{2}$ $O\left(\min\left\{(FG)^{1/3},G^{1/2+\epsilon}M^{1/4}\right\}\right)$ [31, 7] $O((FG)^{1/3})$ [31]
@@
We use Theorem 1 to derive a lower bound for claw detection with a small range. For this, we need a lower bound on the minimum degree of a polynomial that approximates the claw property (for input functions with a possibly large common range). It turns out that for any constant $0<\epsilon<1/2$, the minimum degree of a polynomial that $\epsilon$-approximates the claw property for function pairs $f\colon[F]\to[FG]$ and $g\colon[G]\to[FG]$ is $\Omega\left((FG)^{1/3}\right)$. By combining this with Theorem 1, we can obtain a lower bound on the minimum approximate degree of the claw property for the range size at least $F+G$ of input functions.
@@
Let $F,G\in\mathbb{N}$ be such that $F\leq G$. For any constant $0<\epsilon<1/2$ and any $M\geq F+G$, the minimum degree of a polynomial that $\epsilon$-approximates the claw property for a function pair in $M]^{[F]}\times[M]^{[G]}$ is $\Theta(\sqrt{G}+(FG)^{1/3})$, where the lower bound $\Omega(\sqrt{G})$ holds for all $M\geq 2$.
By the polynomial method, Theorem 3 implies the optimal lower bound on the bounded-error quantum query complexity of both claw detection and claw finding, where the optimality follows from the quantum algorithm [31] that finds a claw.
@@
Let $F,G\in\mathbb{N}$ be such that $F\leq G$. For every $M\geq F+G$, the (query-)optimal quantum algorithm that computes a claw (or detects the existence of a claw) with a constant error probability requires $\Theta(\sqrt{G}+(FG)^{1/3})$ queries for a given pair of functions in $M]^{[F]}\times[M]^{[G]}$. More precisely, the optimal quantum query complexity is $\Theta((FG)^{1/3})$ for every $M\geq F+G$ if $F\leq G\leq F^{2}$ and $\Theta(\sqrt{G})$ for every $M\geq 2$ if $G>F^{2}$.
Ambainis, Balodis and Iraids [7] investigated upper and lower bounds for the claw problem for input function pairs with much smaller ranges. They proved the lower bound $\Omega\left(N^{1/2}M^{1/6}\right)$ when the two input functions has the same domain size $N$ and the range size $M$ with $2\leq M<N$. Their proof is based on the lower bound $\Omega(N^{2/3})$ on the number of quantum queries required to detect a claw for a given pair of functions in $N]^{[N]}\times[N]^{[N]}$, which follows by reducing from element distinctness. However, if the domains of two input functions are different in size, their proof breaks down. By generalizing their proof using Theorem 4, we obtain a lower bound for functions with different-size domains and very small common ranges.
@@
Let $F,G,M\in\mathbb{N}$ be such that $F\leq G\leq F^{2}$ and $M<F+G$. Then, for every $M\in[2,F+G-1]$, the quantum query complexity of detecting the existence of a claw for a given function pair $(f,g)$ in $M]^{[F]}\times[M]^{[G]}$ is lower-bounded by
@@
Note that the condition of $F\leq G\leq F^{2}$ and $M<F+G$ imposed in Theorem 5 handles the parameter range that Theorem 4 does not cover. Setting $F=G=N$ in the theorem recovers the lower bound $\Omega\left(N^{1/2}M^{1/6}\right)$ [7].
To evaluate our bound in Theorem 5, we will compare it with the lower bound obtained by straightforwardly reducing claw detection for two functions with domains of the same size and applying the lower bound given in  [7]. Let us write $\mathsf{claw}_{(a,b)\to c}$ to mean the decision problem of detecting the existence of a claw for given two functions in $c]^{[a]}\times[c]^{[b]}$. The reduction from $\mathsf{claw}_{(G,G)\to M}$ to $\mathsf{claw}_{(F,G)\to M}$ is as follows, where we assume for simplicity that $G$ is divisible by $F$, i.e., $F|G$ (it is easy to extend the reduction to the general case by padding dummy elements to one of the domains). Let $(f,g)$ be a pair of input functions to $\mathsf{claw}_{(G,G)\to M}$. We first partition the domain $[G]$ of $f$ into $G/F$ blocks of size $F$: $B_{1},\dots,B_{G/F}$. For randomly chosen $i\in[G/F]$, we run a $Q$-query quantum algorithm for $\mathsf{claw}_{(F,G)\to M}$ on $f$ restricted to the domain $B_{i}$ and $g$, and apply the quantum amplitude amplification [12, 20] to the whole algorithm. The resulting quantum algorithm makes $O\left(\sqrt{G/F}\cdot Q\right)$ quantum queries. Since the quantum query complexity of $\mathsf{claw}_{(G,G)\to M}$ for $M<G$ is $\Omega(G^{1/2}M^{1/6})$ [7], we have $Q=\Omega(F^{1/2}M^{1/6})$. Together with trivial lower bound $\Omega(\sqrt{G})$, we obtain the lower bound $\Omega(\sqrt{G}+F^{1/2}M^{1/6})$. This bound is weaker than our $\Omega(\sqrt{G}+F^{1/3}G^{1/6}M^{1/6})$. To see this, assume $G=F^{1+\epsilon}$ for $0<\epsilon\leq 1$, satisfying $F<G\leq F^{2}$. In this case, we have
@@
Tab.  1 summarizes the quantum query complexity for claw detection for functions $f\colon[F]\to[M]$ and $g\colon[G]\to[M]$, assuming $F\leq G$ without loss of generality. The upper bound $O(G^{1/2+\epsilon}M^{1/4})$ in the case of $M<F+G$ for $F\leq G\leq F^{2}$ follows from the upper bound [7] for $\mathsf{claw}_{(G,G)\to M}$ (i.e., claw detection for the same domain sizes) via a simple embedding the domain $[F]$ into $[G]$.
The most interesting open problem is finding the tight bound in the case of $M\leq F+G$ for $F\leq G\leq F^{2}$. More precisely, it follows from Tab.  1 that the tight bound $\Theta((FG)^{1/3})$ for $M\geq F+G$ also holds for all $M=\Omega(G)$. On the other hand, if $M=\Theta(1)$, then $\Omega(\sqrt{G})$ is tight, since Grover’s search algorithm achieves $O(\sqrt{G})$ queries. Thus, the tight bound is open when $M$ is asymptotically smaller than $G$ and larger than $2$.
@@ 1.4 Technical Outline
The starting point of the proofs of Theorems 1 and 2 is the proof idea for the single-symmetric properties introduced by Ambainis [4] (we hereafter refer to "single-symmetric" just as "symmetric"). Let $\Phi$ and $\Phi^{\prime}$ be the symmetric Boolean functions defined in Sec.   1.2. Recall that $\Phi$ and $\Phi^{\prime}$ are functions in $\{x_{ij}\colon(i,j)\in[n]\times[m]\}$ and $\{x_{ij}\colon(i,j)\in[n]\times[n]\}$, respectively. In addition, define $z_{j}$ as the number of preimages of $j$ via $f$, i.e., $z_{j}=\left|f^{-1}\left(j\right)\right|$. Then, Ambainis’s proof consists of two claims. The first claim is that the following are equivalent:
@@ 2. $i$
These claims together imply that the minimum degree of a polynomial that approximates $\Phi$ is at most $d$, if and only if the minimum degree of a polynomial approximating $\Phi^{\prime}$ is at most $d$.
We want to generalize these two claims to the bisymmetric case (the $k$-symmetric case is a straightforward extension of the bisymmetric case). Recall in the bisymmetric case that $f\colon[F]\to[M]$ is identified with $FM$ Boolean variables $\{x_{ij}:i\in\left[F\right],j\in\left[M\right]\}$, and $g\colon[G]\to[M]$ is identified with $GM$ Boolean variables $\{y_{kj}:k\in\left[G\right],j\in\left[M\right]\}$. In addition, we define $z_{j}$ as the number of preimages of $j$ via $f$, i.e., $z_{j}=\left|f^{-1}\left(j\right)\right|$, and define $w_{j}$ as the number of preimages of $j$ via $g$, i.e., $w_{j}=\left|g^{-1}\left(j\right)\right|$. Let $\Phi$ and $\Phi^{\prime}$ be the bisymmetric Boolean function in $\{x_{ij}\}$ and $\{y_{kj}\}$ as defined in Sec.   1.3. The first claim in the bisymmetric case is that the following are equivalent:
@@
For the first claim in the symmetric case, it is easy to show that item $1$ implies item $2$ by using the trivial relation $z_{j}=\sum_{i}x_{ij}$. This also works in the bisymmetric case. To show the converse in the symmetric case, we consider the average $\mathsf{E}_{\pi}[P]$ of $P$ over all permutations $\pi$ over the domains of $f$, since $\mathsf{E}_{\pi}[P]$ also approximates $\Phi$ due to the symmetry of $\Phi$ over the domains. Since we are interested in the degree of the resulting polynomial, it suffices to examine the average of each product term $x_{i_{1}j_{1}}\cdots x_{i_{k}j_{k}}$ in $P$, and it turns out that $\mathsf{E}_{\pi}\left[x_{i_{1}j_{1}}\cdots x_{i_{k}j_{k}}\right]$ is a polynomial in $z_{j}$’s of degree at most $d$. In the bisymmetric case, we take the average $\mathsf{E}_{\pi_{F}\times\pi_{G}}[P]$ of $P$ over all permutations $\pi_{F}$ over the domains of $f$ and all permutations $\pi_{G}$ over the domains of $g$. Note that each product term in $P$ is of the form of $x_{i_{1}j_{1}}\cdots x_{i_{m}j_{m}}y_{k_{1}j_{1}}\cdots y_{k_{l}j_{l}}$. To obtain the average of each term, we apply permutations $\pi_{F}$ and $\pi_{G}$ independently, and thus it holds that
@@
For the second claim, it is more involved to generalize the symmetric case to the bisymmetric case since the permutation over the range is common to the two functions. It is easy to see that item (i) implies (ii) in the symmetric case since we can obtain $Q^{\prime}$ by setting $z_{n+1}=\cdots=z_{m}=0$ in $Q$. This also works in the bisymmetric case. For the converse, let us first review the symmetric case. Since every $z:=(z_{1},\dots,z_{m})$ corresponding to a function $f:\left[n\right]\rightarrow\left[m\right]$ has at most $n$ nonzero coordinates, there exists a permutation $\sigma$ over $[m]$ that moves all non-zero coordinates to the first $n$ coordinates, that is, $\sigma z=\left(z_{\sigma^{-1}\left(1\right)},\ldots,z_{\sigma^{-1}\left(m\right)}\right)$ satisfies that $z_{\sigma^{-1}\left(j\right)}=0$ for every $\sigma^{-1}\left(j\right)\geq n+1$. This action of $\sigma$ does not change the value of $\Phi$: $\Phi\left(f\right)=\Phi\left(\sigma f\right)$. Since $\sigma f$ can be regarded as a function in $\left[n\right]^{\left[n\right]}$, it must hold that $\Phi\left(\sigma f\right)=\Phi^{\prime}\left(\sigma f\right)$. By the assumption, $\Phi^{\prime}\left(\sigma f\right)$ is approximated by $Q^{\prime}\left(\sigma z\mid_{n}\right)$, where $\sigma z\mid_{n}$ is the first $n$ variables of $\sigma z$, i.e., $z_{\sigma^{-1}\left(1\right)},\ldots,z_{\sigma^{-1}\left(n\right)}$. We then define from $Q^{\prime}$ a symmetric polynomial $Q$ in $m$ variables, $z_{1},\ldots,z_{m}$, such that, for $z$ corresponding to an arbitrary $f$,
@@
where $\sigma$ is a particular permutation defined above depending on $f$. This implies that $Q\left(z\right)$ approximates $\Phi\left(f\right)$, since $Q^{\prime}\left(\sigma z\mid_{n}\right)$ approximates $\Phi^{\prime}\left(\sigma f\right)=\Phi\left(\sigma f\right)=\Phi\left(f\right)$. Since we can assume without loss of generality that $Q^{\prime}$ is a symmetric polynomial, the fundamental theorem of symmetric polynomials implies that $Q^{\prime}$ is represented as a polynomial in the elementary polynomials over $z_{1},\dots,z_{n}$. We then replace these elementary polynomials in $Q^{\prime}$ with the corresponding polynomials over $z_{1},\dots,z_{m}$ to obtain $Q$.
In the bisymmetric case, $Q$ is a polynomial in two sets of variables: $z_{1},\dots,z_{M}$ and $w_{1},\dots,w_{M}$; $Q^{\prime}$ is a polynomial in two sets of variables $z_{1},\dots,z_{M^{\prime}}$ and $w_{1},\dots,w_{M^{\prime}}$ such that $F+G\leq M^{\prime}<M$, since there exists a permutation $\sigma$ over $[M]$ that moves all non-zero coordinates in $z_{1},\dots,z_{M}$ and $w_{1},\dots,w_{M}$ to the first $M^{\prime}$ coordinates, respectively. To obtain $Q$ from $Q^{\prime}$, we utilize a certain symmetry of $Q^{\prime}$. However, $Q^{\prime}$ is not invariant under the action of permutations over $2M^{\prime}]$ to $2M^{\prime}$ variables $z_{1},\dots,z_{M^{\prime}},w_{1},\dots,w_{M^{\prime}}$, but it is invariant under the action of $\tau\times\tau$ for permutation $\tau$ over $M^{\prime}]$ such that $(z_{1},\dots,z_{M})\mapsto\left(z_{\tau^{-1}\left(1\right)},\ldots,z_{\tau^{-1}\left(M^{\prime}\right)}\right)$ and $(w_{1},\dots,w_{M})\mapsto\left(w_{\tau^{-1}\left(1\right)},\ldots,w_{\tau^{-1}\left(M^{\prime}\right)}\right)$. Here, the notion of multisymmetric polynomials comes in. Roughly speaking, multisymmetric polynomials are those invariant under the action of permutations to the set of vectors of variables, instead of variables. In our case, let us consider the set of vectors $(z_{1},w_{1}),\dots,(z_{M^{\prime}},w_{M^{\prime}})$, each of which consists of two variables. Then, $Q^{\prime}$ is invariant under the action of the permutation $\tau$ on this set of vectors. Now that we have entered the world of multisymmetric polynomials, we can use the fundamental theorem of those polynomials. More concretely, $Q^{\prime}$ is represented as a polynomial in power-sum polynomials (defined in Sec.   2.4) in $(z_{1},w_{1}),\dots,(z_{M^{\prime}},w_{M^{\prime}})$. Then, we can obtain $Q$ by replacing each occurrence of power-sum polynomials in $Q^{\prime}$ with the corresponding polynomials in $(z_{1},w_{1}),\dots,(z_{M},w_{M})$. This approach is easily generalized to the $k$-symmetric case. We just consider the symmetry under the action of permutations to the set of vectors $\left(z_{1}^{(1)},\dots,z_{1}^{(k)}\right),\dots,\left(z_{M^{\prime}}^{(1)},\dots,z_{M^{\prime}}^{(k)}\right)$, where $z_{i}^{(\ell)}=\left\lvert f_{\ell}(j)\right\rvert$.
@@
To provide a lower bound for the claw problem with a small range using Theorem 1, we need a lower bound on the minimum degree of a polynomial that approximates the claw property for some common range size of input functions. For this, we use Zhang’s reduction [33] from the collision problem to the claw problem. Suppose that there exists a polynomial $P^{\prime}$ of degree $d$ that 0.1-approximates the claw property for functions $f\in[FG]^{[F]}$ and $g\in[FG]^{[G]}$ such that $F\leq G$. Based on $P^{\prime}$, we will then construct a polynomial $P$ of degree $d$ that (41/86)-approximates the collision property for functions $h\in[M]^{[M]}$ for $M=FG$ with the promise that $h$ is either one-to-one or two-to-one. Here we mean by the collision property the Boolean function $\Phi$ which is $\mathsf{true}$ if and only if the input function $h$ is two-to-one. Since the minimum degree of the latter polynomial is $\Omega(M^{1/3})$ [2, 4], we obtain $d=\Omega((FG)^{1/3})$. Theorem 1 with this bound implies Theorem 3, from which we obtain Theorem 4 with the polynomial method.
To obtain polynomial $P$ from $P^{\prime}$, let $P^{\prime}_{ST}$ for disjoint subsets $S,T$ of domain $[M]$ of $h$ be a polynomial of degree $d$ (in variables $x_{ij}$ for $i\in S\sqcup T$ and $j\in[M]$) that $\epsilon$-approximates the claw property on function pairs $M]^{S}\times[M]^{T}$ to $\{0,1\}$. Then, define a polynomial $P$ as the average of $P^{\prime}_{ST}$ over uniformly random disjoint subsets $S,T$ of size $F$ and $G$, respectively: